OpenStudy (anonymous):

Use one of your formulas to find the degree equivalent of . What is the result?

OpenStudy (jack1):

... not enough data... is there more?

OpenStudy (anonymous):

Use one of your formulas to find the degree equivalent of 11pie/6 . What is the result?

OpenStudy (anonymous):

not sure what my formulas are or how to solve it

OpenStudy (anonymous):



OpenStudy (jack1):

\(\Large 11 \frac{\pi}6\) in degrees, yeah? ok, where u see PI... put 180 this will solve mathematically (as pi radians = 180 degrees)

OpenStudy (jack1):

so 11 x 180 / 6 = ???

OpenStudy (anonymous):

330 ?

OpenStudy (jack1):

perfect @jaesierra

OpenStudy (anonymous):

thanks :)

OpenStudy (jack1):

np ;)